TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes


Created by Main.MircoCoccoli on 2006-11-27 - 19:08

RAT - Preparation for Cool Down sec.78 - Nov 27

PDF RAT_20061127.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x27x19x02/RAT_20061127.pdf

Main.MircoCoccoli - 27 Nov 2006 - 19:02

Open Issues:

AC non-conformity

  • 21.11.06 Transport and installation of the heater system of the current leads for RR77 and UA83 (P.Chambouvet: about 2 weeks of work starting Mon 27 Nov)
  • 21.11.06 Non-conformity on instrumentation cables (temperature sensors on current leads) (A.Suraci: 4 cables are in repair)
  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converters & VAC.
DC cabling
  • 25.11.06 Functional test on the leads heaters: DFBMA, DFBMC, DFBX
  • 25.11.06 120A and 600A positioning on DFBAN and DFBMH
  • 23.11.06 Cable temporary positioning on the transport side of the tunnel (A.Vergara: next year)
  • 22.11.06 Pre-connection of 120A cables in LSSR7
  • 20.11.06 Galvanic insulation installation on all DFBs
  • 13.11.06 Non-conformity of the DC cabling of the orbit corrector power converters
  • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA and DFBMC
Assembly - Vacuum - Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO
  • 22.11.06 Control of quench valve
  • 20.11.06 DFB instrumentation cables to be connected (D.Bozzini: 80% done)
  • 17.11.06 Instrumentation cable HV tests in LSS8L (A.Suraci: 4 cables to be tested)
DFB Commissioning
  • 13.11.06 ElQA-TP4-A for all DFBs between 7 and 8 (D.Bozzini: ongoing according to schedule (next: DFBMA, DFBMC, DFBX); in week 48 will begin ElQA-TP4-B, i.e. HV tests)
  • 17.11.06 Pumping of insulation vacuum for all DFBs between 7 and 8 (P.Chambouvet: DFBAN done; next: DFBMA, DFBMC, DFBX)
Safety General remark: written communication in advance to announce operations!!!
  • 23.11.06 Wooden structure for UA access restriction (that will allow transport).
  • 23.11.06 Preparation of a prototype insulating protection (derived from a plastic tube) of the monorail by J.C.Guillaume and approval by J.Etheridge.
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (33L8).
  • 23.11.06 Planning of consignations/deconsignations
  • 16.11.06 Cool down safety procedure & access conditions
Preparation of powering test
  • 22.11.06 Pretest of 60A converters in sector 78, location C16L8. (B.Bellesia: stopped due to LQA hardware problems; expected end by Tuesday 28 Nov)
  • 22.11.06 Installation of flexible cables by CV (UA83)
  • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be checked.
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7
  • 13.11.06 Quench protection continuity tests and cables assignment LLS7R (DFBMH, DFBAN) and LSS8L (DFBMC, DFBMA and DFBX)

Edit, attach file or add comment


Created by Main.MirkoPojer on 2006-11-27 - 11:34

RAT at Point 5, Nov 27

PDF RAT_20061127.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x11x27x11x29/RAT_20061127.pdf

Comments:

Open issues

  • 23.11 Ripples in RQT12R5B1 power converter.
  • 23.11 Cables inversion on rack RYMCA04.
  • 20.11 Problem with water conductivity in DQS: reason known, insulation test still to be performed.
  • 13.11 DC cables high voltage test (RR57 done. Waiting UJ56)
  • 13.11 Test de circulation ED (waiting UJ56)
  • 22.11 Full power ventilation test in RR57 – Friday 8th Dec. 8:30-12:00
  • 22.11 Fire detection – awaiting certification

Edit, attach file or add comment

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2006-11-01 - JacekSzkutnik
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login